# Optimal Virtual Model Control for Robotics: Design and Tuning of Passivity-Based Controllers

### v2

@@ V-A Optimal PD control of a cart
where $q$ and $\dot{q}$ are position and velocity of the cart, $d$ is an external force acting on the cart, and $n_{q}$ and $n_{\dot{q}}$ are disturbances affecting position and velocity measures. $m$ is the mass of the cart and $k$ and $b$ are the proportional and derivative parameters of the controller, which correspond to virtual mechanical stiffness and damping, respectively. These are collected in the parameter vector ${\bm{\theta}}$.
@@
with $\eta_{i}$ randomly sampled from the space of 3D unit vectors. Frequencies are sampled exponentially between $0.1$rad/s and $500$rad/s. The lower frequencies are sampled more densely, as this is where the dynamics of interest occur.
Figure 11 shows a comparison between the true gain $\mathcal{L}_{2}$ obtained by $H_{\infty}$ synthesis [48], and the result of our optimization, as the size of the set of scenarios grows. As a greater number of frequencies are sampled, the cost landscape more closely resembles the ‘true’ cost. The white lines represent the trajectory of the parameters $k$ and $b$ as the optimization converges [^1] We have used the ADAM optimizer with an initial step size of $1.0$.
@@
The adversarial approach provides an alternative to dense sampling, dramatically reducing the number of scenarios. Figure 12 shows the optimization based on the adversarial approach. The two steps of optimization (15a) and (15b) run on scenarios parameterized four elementsvectors $\omega^{i}=[\,\omega_{1}^{i},\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,]^{T}$, specifically
$\Omega_{i}=\omega_{1}^{i}\quad\eta^{T}=[\,\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,],$ (22)
where $\omega_{1}^{i}$ is sampled in the range $[0,500]$222…by taking the exponential of a uniformly sampled parameter in the range $[0,\log(500)]$, so that there is some bias towards lower frequencies., and $[\,\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,]$ is sampled from the unit normal distribution. An adversarial optimizer is initialized [^3] An ADAM optimizer with initial step size $2.0$ and the ‘worst case’ probing signal is obtained by approximating (15b) via 500 iterations (the controller does not change in this phase). Then, the controller adjustment (15a) is resumed. And so on. Figure 12 illustrates the cost function for increasing scenario sizes, starting from a very small setting of three signals. Each new scenario is optimally chosen by (15b).
The summary in Table I and the visual comparison between Figure 12 and Figure 11 show that the adversarial case is in closer agreement with the true cost landscape (top left of Figure 11). Remarkably, this is achieved with only 22 scenarios, a lot less than the 400 scenarios of the sampling case. As the computational cost scales linearly with the number of scenarios, the adversarial approach offers an order-of-magnitude improvement in the tuning time.
@@
TABLE I: Controlled cart parameters and estimated $\mathcal{L}_{2}$ cost.
@@
in (13). In this case, we use directly adversarial optimization. For simplicity, we use the parameterization (21) as in the $\mathcal{L}_{2}$ case, even if sinusodial signals are not necessarily the best choice to calculate the $\mathcal{L}_{\infty}$ cost. The results are illustrated in Figure 13, which shows significant similarities and differences compared to the cost landscape of Figure 12. The optimal proportional gain $k$ is smaller, while the derivative gain $b$ remains unchanged between the two cases. In contrast to the $\mathcal{L}_{2}$ case, we cannot offer here a comparison with the ‘true’ $\mathcal{L}_{\infty}$ to $\mathcal{L}_{\infty}$ gain, as this control synthesis problem is still unsolved, even in the linear setting.
V-B Optimal Virtual Model Control for Laparoscopic Surgery
We demonstrate the applicability of the optimization approach to useful robotic systems, which have several degrees of freedom, and a more complex virtual-mechanism controller, by applying the tuning procedure to the virtual instrument controller for surgery described in III-B, with an additional virtual joint damper on J1 and on J4 to arrest null space motion in the simulation/tuning (which features no modelled joint-damping).
The robot manipulator kinematics is that of Franka Emika Research 3. We use the inertial parameters determined in [51]. The virtual instrument is given an inertia of $0.05I\,\mathrm{kg\,m}^{2}$ and an inertance of $1.0\,\mathrm{kg}$ by attaching a linear between its centre and ground [^4] This is equivalent to endowing a point mass of 1kg in zero gravity.. For simplicity, these parameters are set and will not be tuned in this paper. There are $3$ virtual springs and $5$ virtual dampers to optimize, which leads to a vector of control parameters ${\bm{\theta}}$ of eight elements.
@@ V-B Optimal Virtual Model Control for Laparoscopic Surgery
to limit the maximum stiffness and damping values. We penalize components of ${\bm{\theta}}$ above $\ln{}(3000)$, as robot actuators cannot realize excessive stiffness / damping. This corresponds to a soft limit on the stiffnesses at $3000$ and the damping coefficients at 30.030.030.030.0555This is a crude way to avoid reaching the performance limits of our robot. A better identification of the bandwidth limits of the joints is needed to develop a more principled approach, beyond the scope of this paper.. The corresponding values pre and post optimization are shown in Table II.
@@
EE-ref stiffness     $6.91$ $1000$     $8.02$ $3056.4$
Base-VI stiffness     $6.91$ $1000$     $7.22$ $1368.4$
EE-VI stiffness     $6.91$ $1000$     $8.00$ $2993.7$
EE-ref damping     $6.91$ $1000$     $8.01$ $30.1$
Base-VI damping     $6.91$ $1000$     $7.86$ $25.8$
EE-VI damping     $6.91$ $1000$     $7.97$ $28.8$
J1 damping     $6.91$ $1000$     $3.74$ $0.421$
J4 damping     $6.91$ $1000$     $6.27$ $5.29$
TABLE II: Initial and tuned parameters (rounded)
where $W_{{\bm{d}}_{ee}}=8I$, $W_{{\bm{d}}_{rcm}}=4I$, $W_{{\bm{n}}_{q}}=0.002I$, $W_{{\bm{n}}_{\dot{q}}}=0.1I$, and $W_{{\bm{r}}}=0.05I$. We assume that the input ${\bm{w}}$ is a vector of sinusoids with maximal norm of $1$, therefore the disturbance forces peak at $8$N on the end-effector or $4$N on the port, and the reference requires a displacement not larger than $5$ cm. Noise on the velocity variables is greater than on the position variables, as velocities are ordinarily obtained by numerical differentiation, which amplifies noise.
The performance output ${\bm{y}}$ captures relevant position errors
${\bm{y}}=Q\begin{bmatrix}{\bm{x}}_{ee}-{\bm{r}}_{ee}\\ {\bm{e}}_{rcm}\\ \end{bmatrix},$ (26)
where ${\bm{x}}_{ee}$ denotes the position of the end effector and ${\bm{e}}_{rcm}$ represents the ‘port error’, that is, the minimal displacement between the centre of the port and the longitudinal axis passing through the instrument, as shown in Figure 14. The output weighting matrix is chosen to emphasize performance at the end-effector, for which precision is more important.
$Q^{2}=\begin{bmatrix}I_{3\times 3}&0\\ 0&0.2I_{2\times 2}\end{bmatrix}$ (27)
Figure 14: Performance outputs considered for the optimization of the parameters of the virtual instrument. $e_{rcm}$ is a length 2 vector representing the x/y position of the RCM in the instrument frame. The instrument axis is along the z axis of the instrument frame, thus $\parallel e_{rcm}\parallel_{2}$ is the shortest distance from the instrument axis to the RCM. $x_{ee}-r_{ee}$ is the vector from the reference to the end-effector, and thus $\parallel x_{ee}-r_{ee}\parallel_{2}$ is the distance from the end-effector to the reference.
For the numerical optimization, we initialize the robot simulations at a rest position, with the end effector near the desired position and the instrument passing through the port. We optimise over $7$ scenarios, summarized in Table III. Each scenario includes a combination of noise ${\bm{n}}=[{\bm{n}}_{q}^{T},{\bm{n}}_{\dot{q}}^{T}]^{T}$, disturbance ${\bm{d}}=[{\bm{d}}_{ee}^{T},{\bm{d}}_{rcm}^{T}]^{T}$, and reference signal ${\bm{r}}$, which are otherwise zero. We consider
@@
TABLE III: Exogenous Signals present in each Scenario
We test the optimized parameters on each scenario, as illustrated in Figure 15, right. The exogenous inputs included in each scenario are indicated in Table III, alongside the (rounded) tuned parameters. The comparison between the initial parameters and optimal tuning shows a consistent improvements in the cost metric in response to disturbances and references. The case scenario 1 shows how the optimizing over all scenarios may increase the gain in a specific scenario. Plots of the achieved performance for Scenarios 2 and 4 are shown in Figure 16. These show the resulting decrease in error at the end-effector and at the RCM in response to the pure disturbance and pure reference signals respectively.
@@
VI Experimental validation
@@
The tip/RCM errors shown in Figure 17 E and F vary depending on the port/pose, but typically remains under 8mm. The RCM error typically remains under 5mm. This reference tracking could potentially be improved by increasing the limit of stiffness; or by the compensation of friction and coriolis/centrifugal forces; or by shaping the inertia matrix of the robot to reduce its apparent inertia. These additional control terms are non-passive, but proven methods if an accurate robot dynamic model is available, and are entirely compatible with the virtual mechanism approach.
Because of our approach and the structure of the controller, particularly its passivity, we expect that the interconnection of the virtual model controller and the real robot should remain stable. This is despite mismatches between the robot model used for the tuning and the robot itself, such as uncertainty in the inertia parameters and the lack of any joint-friction model. Due to the constraint of passing through the ports of the surgical phantom, all of the 4 poses tested, represented in Figure 17 C, deviate from the training pose. This results in a different behaviour to that experienced in simulation and tuning. However, due to the structure of the controller, we still expect reasonable performances. Notably, if the sensing/communication/actuation loop accurately renders the desired torques, the passivity of the controller guarantees that the closed loop system will be stable even if the dynamics of the real robot do not match the simulated dynamics and under interaction with external passive systems, such as the interaction between the instrument and the port. However, a loss of performance is expected due the mismatch between ideal and real parameters of the robot.
@@
Figure 17: Experimental performance of the controller while passing through 4 different ports on a laparoscopy training phantom. Video frames showing port A at t=0s, 1s, …5s, from the side A) and top B) of the phantom respectively. C) Initial poses for of the robot/virtual instrument. D) Tip error x/y trajectories. E) Tip error magnitude. F) RCM error magnitudes. All positions are computed using the kinematics of the robot and measured joint angles.
VII Conclusions

### v3

@@ VI-A Optimal PD control of a cart
where $q$ and $\dot{q}$ are position and velocity of the cart, $d$ is an external force acting on the cart, and ${\bm{n}}_{q}$ and ${\bm{n}}_{\dot{q}}$ are disturbances affecting position and velocity measures. $m$ is the mass of the cart and $k$ and $b$ are the proportional and derivative parameters of the controller, which correspond to virtual mechanical stiffness and damping, respectively. These are collected in the parameter vector $\bm{\theta}$}.
@@
with $\eta_{i}$ randomly sampled from the space of 3D unit vectors. Frequencies are sampled exponentially between 0.10.1rad/s and 500500rad/s. The lower frequencies are sampled more densely, as this is where the dynamics of interest occur.
Fig. 12 shows a comparison between the true gain $\mathcal{L}_{2}$ obtained by $H_{\infty}$ synthesis [Zhou1996], and the result of our optimization, as the size of the set of scenarios grows. As a greater number of frequencies are sampled, the cost landscape more closely resembles the ‘true’ cost. The white lines represent the trajectory of the parameters $k$ and $b$ as the optimization converges [^3] We have used the ADAM optimizer with an initial step size of $1.0$.
@@
The adversarial approach provides an alternative to dense sampling, dramatically reducing the number of scenarios. Fig. 13 shows the optimization based on the adversarial approach. The two steps of optimization (21a) and (21b) run on scenarios parameterized by four elements $\omega^{i}=[\,\omega_{1}^{i},\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,]^{T}$, specifically
$\Omega_{i}=\omega_{1}^{i}\quad\eta^{i}=[\,\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,]^{T},$ (28)
where $\omega_{1}^{i}$ is sampled in the range [0,500][0,500]444…by taking the exponential of a uniformly sampled parameter in the range $0,\log(500)]$, so that there is some bias towards lower frequencies., and $\omega_{2}^{i},\omega_{3}^{i},\omega_{4}^{i}\,]$ is sampled from the standard normal distribution. An adversarial optimizer is initialized [^5] An ADAM optimizer with initial step size $2.0$ and the ‘worst case’ probing signal is obtained by approximating (21b) via 500 iterations (the controller does not change in this phase). Then, the controller adjustment (21a) is resumed. And so on. Fig. 13 illustrates the cost function for increasing scenario sizes, starting from a very small setting of three signals. Each new scenario is optimally chosen by (21b).
The summary in Table I and the visual comparison between Fig. 13 and Fig. 12 show that the adversarial case is in closer agreement with the true cost landscape (top left of Fig. 12). Remarkably, this is achieved with only 22 scenarios, a lot less than the 400 scenarios of the sampling case. As the computational cost scales linearly with the number of scenarios, the adversarial approach offers an order-of-magnitude improvement in the tuning time.
TABLE I: Controlled cart parameters and estimated $\mathcal{L}_{2}$ cost.
@@
in (19). In this case, we use directly adversarial optimization. For simplicity, we use the parameterization (27) as in the $\mathcal{L}_{2}$ case, even if sinusodial signals are not necessarily the best choice to calculate the $\mathcal{L}_{\infty}$ cost. The results are illustrated in Fig. 14, which shows significant similarities and differences compared to the cost landscape of Fig. 13. The optimal proportional gain $k$ is smaller, while the derivative gain $b$ remains unchanged between the two cases. In contrast to the $\mathcal{L}_{2}$ case, we cannot offer here a comparison with the ‘true’ $\mathcal{L}_{\infty}$ to $\mathcal{L}_{\infty}$ gain, as this control synthesis problem is still unsolved, even in the linear setting.
VI-B Optimal Virtual Model Control for Laparoscopic Surgery
We demonstrate the applicability of the optimization approach to useful robotic systems, which have several degrees of freedom, and a more complex virtual-mechanism controller, by applying the tuning procedure to the virtual instrument controller for surgery described in IV-B, with an additional virtual joint damper on J1 and on J4 (i.e. an additional torque proportional to $\dot{q}_{1}$ applied by $u_{1}$, and on J4) to arrest null space motion in the simulation/tuning (which features no modelled joint-damping).
The robot manipulator kinematics is that of Franka Emika Research 3. We use the inertial parameters determined in [Gaz2019]. The virtual instrument is given an inertia of $0.05I\,\mathrm{kg\,m}^{2}$ and an inertance of $1.0\,\mathrm{kg}$ by attaching a linear inerter between its centre and ground [^6] This is equivalent to endowing a point mass of 1kg in zero gravity.. For simplicity, these parameters are set and will not be tuned in this paper. There are $3$ virtual springs and $5$ virtual dampers to optimize, which leads to a vector of control parameters $\bm{\theta}$} of eight elements.
@@ VI-B Optimal Virtual Model Control for Laparoscopic Surgery
to limit the maximum stiffness and damping values. We penalize components of $\bm{\theta}$} above $\ln{}(3000)$, as robot actuators cannot realize excessive stiffness / damping. This corresponds to a soft limit on the stiffnesses at $3000$ and the damping coefficients at 30.030.0777This is a crude way to avoid reaching the performance limits of our robot. A better identification of the bandwidth limits of the joints is needed to develop a more principled approach, beyond the scope of this paper.. The corresponding values pre and post optimization are shown in Table II. The exact placement of the springs is shown in Fig. 1. The following abbreviations are used: End-Effector $\rightarrow$ EE, Virtual Instrument $\rightarrow$ VI, reference $\rightarrow$ ref, Joint 1 $\rightarrow$ J1.
TABLE II: Initial and tuned parameters (rounded).
@@
[…]
where $W_{{\bm{d}}_{ee}}=8I$, $W_{{\bm{d}}_{rcm}}=4I$, $W_{{\bm{n}}_{q}}=0.002I$, $W_{{\bm{n}}_{\dot{q}}}=0.1I$, and $W_{{\bm{r}}}=0.05I$. We assume that the input $\bm{w}$} is a vector of sinusoids with maximal norm of $1$, therefore the disturbance forces peak at 88N on the end-effector or 44N at the RCM, and the reference requires a displacement not larger than $5$ cm. In this tuning setup, the weights are chosen to normalise quantities of different units and magnitudes, ensuring that reference motions, disturbances and noise signals are of a reasonable scale. Noise on the velocity variables is greater than on the position variables, as velocities are ordinarily obtained by numerical differentiation, which amplifies noise.
The performance output $\bm{y}$} captures relevant position errors
${\bm{y}}=Q\begin{bmatrix}{\bm{x}}_{ee}-{\bm{r}}_{ee}\\ {\bm{e}}_{rcm}\\ \end{bmatrix}$, (32)
where ${\bm{x}}_{ee}$ denotes the position of the end effector and ${\bm{e}}_{rcm}$ represents the ‘RCM error’, that is, the minimal displacement between the RCM and the longitudinal axis passing through the instrument, as shown in Fig. 15. The output weighting matrix is chosen to emphasize performance at the end-effector, for which precision is more important.
$Q^{2}=\begin{bmatrix}I_{3\times 3}&0\\ 0&0.2I_{2\times 2}\end{bmatrix}$ (33)
Figure 15: Performance outputs considered for the optimization of the parameters of the virtual instrument. $e_{rcm}$ is a length 2 vector representing the x-y position of the RCM in the instrument frame. The instrument axis is along the z axis of the instrument frame, thus $\parallel e_{rcm}\parallel_{2}$ is the shortest distance from the instrument axis to the RCM. $x_{ee}-r_{ee}$ is the vector from the reference to the end-effector, and thus $\parallel x_{ee}-r_{ee}\parallel_{2}$ is the distance from the end-effector to the reference.
For the numerical optimization, we initialize the robot simulations at a rest position, with the end effector near the desired position and the instrument passing through the RCM. We optimise over $7$ scenarios, summarized in Table III. Each scenario includes a combination of noise ${\bm{n}}=[{\bm{n}}_{q}^{T},{\bm{n}}_{\dot{q}}^{T}]^{T}$, disturbance ${\bm{d}}=[{\bm{d}}_{ee}^{T},{\bm{d}}_{rcm}^{T}]^{T}$, and reference signal $\bm{r}$}, which are otherwise zero. We consider
${\bm{n}}_{q}$ $={\bm{n}}_{\dot{q}}=[1\,1\,1\,1\,1\,1\,1]^{T}\frac{\sin(50t)}{\sqrt{7}}$ (34a)
${\bm{d}}_{ee}$ $={\bm{d}}_{rcm}=[1\,1\,1]^{T}\frac{\sin(2t)}{\sqrt{3}}$ (34b)
${\bm{r}}$ $=[1\,0\,0]^{T}\frac{\sin(3t)}{\sqrt{3}}.$ (34c)
@@
As indicated by the cost function (19), we optimize one set of parameters by minimising the max cost over all seven scenarios simultaneously, resulting in a single controller that gives the “best-worst” performance over the seven scenarios.
@@
TABLE III: Exogenous Signals present in each Scenario
@@
We test the optimized parameters on each scenario, as illustrated in Fig. 16, right. The exogenous inputs included in each scenario are indicated in Table III, alongside the (rounded) tuned parameters. The comparison between the initial parameters and optimal tuning shows a consistent improvements in the cost metric in response to disturbances and references. The case scenario 1 shows how the optimizing over all scenarios may increase the gain in a specific scenario. Plots of the achieved performance for Scenarios 2 and 4 are shown in Fig. 17. These show the resulting decrease in error at the end-effector and at the RCM in response to the pure disturbance and pure reference signals respectively.
@@
VI-C Scalability of the Optimal Laparoscopic Controller
The small set of selected scenarios in the previous section allows for a detailed comparison of the achieved performance, summarized by the bar chart plot of Figure 16. These examples are chosen for clarity, illustrating the method on a few representative signals that can be easily explained and compared. In this section, we demonstrate that the tuning technique generalises to a larger number of scenarios and more complex signals.
In what follows, scenarios are generated by randomly by sampling RCM positions, noise signals, disturbance signals, and reference signals. For each scenario, the RCM position is drawn from a Gaussian distribution centred on the nominal position (as in the previous Section), with a standard-deviation of $0.05\text{\,}\mathrm{m}$. Each noise, disturbance, and reference signal is constructed as a sum of three sinusoids in a random direction of the form $\hat{n}_{u}\sum_{i=1}^{3}A_{i}\sin(\omega_{i}t)$. The frequencies $\omega_{i}$ are independently sampled from uniform distributions within the ranges shown in Table IV. The amplitudes $A_{i}$ are sampled in the range $[0,1.0]$ and scaled by $W_{w}$, as defined in (31). The direction vector $\hat{n}_{u}$ is obtained by sampling the appropriate number of values from a unit normal distribution, assembling them into a vector, and normalizing.
Each scenario begins with the robot passing through the RCM, and the target tip position initially offset by $100\text{\,}\mathrm{m}\mathrm{m}$ in the $x$-direction and $100\text{\,}\mathrm{m}\mathrm{m}$ in the $z$-direction. By varying the RCM position and thus the robot pose, a wider portion of the system’s nonlinear dynamics is explored. The randomised directions, amplitudes, and frequencies of the noise, disturbance, and reference signals further enrich the excitation space. This demonstrates the flexibility of the approach: there is nothing special about the chosen class of signals. Any desired input signals, representative of the conditions under which performance is to be optimised, can be used.
[…]
The results in Fig. 18 and final parameters in Table V show a higher overall cost, reflecting the increased challenge of the randomised signals. However, the optimised parameters follow similar trends—high stiffness and damping near the regularisation limits, with the notable exception of the EE-VI stiffness being much lower.
We remark that the goal of this section is to demonstrate the scalability of the propose optimization procedure, rather than to produce a complete control design for laparoscopic surgery. The latter is beyond the scope of the current paper and would require a high-fidelity system model and carefully constructed excitation signals, representative of key operating conditions.
Figure 18: Training results on the larger set of scenarios [Left] Log-log cost vs training step. [Right] Initial and final costs, $\frac{\parallel{\bm{y}}\parallel_{2}}{\parallel{\bm{w}}\parallel_{2}}$, on each individual scenario.
TABLE IV: Frequency ranges for the extended set of excitation signals.
    Freq. range (rad/s)
Signal     Min Max
${\bm{n}}_{q}$     $25.0$ $100.0$
${\bm{n}}_{\dot{q}}$     $25.0$ $100.0$
${\bm{d}}_{ee}$     $0.0$ $5.0$
${\bm{d}}_{rcm}$     $0.0$ $5.0$
$\bm{r}$}     $0.0$ $5.0$
VII Experimental validation
We validate the optimal controller of Section VI-B experimentally on a Franka Emika manipulator. The robot is illustrated in Fig. 1, and it is tested passing through 4 different ‘incisions‘ on the surgical phantom of Fig. 7, which are openings in the plastic body with rubber aperture. The RCM location is determined by moving the robot so that the tip of the instrument is at the centre of the aperture, and measuring the location. Then the robot is moved by hand so that the instrument is inserted into the phantom, and the controller is started. To mimic a surgical reference, we apply a reference signal around the starting tip location of $\bm{r}=[0.02\sin(t),0.01\sin(2t),0]^{T}$, a figure-8 motion in the x-y plane with peak-peak amplitude 40mm, with frequency components at $1\text{\,}\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$ and $2\text{\,}\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$ as shown in Fig. 19(d).
The controller runs on a PC that communicates with the robot via a C++ program based on the “libfranka” library. The virtual model controller is written in Julia using the VMRobotControl.jl library [VMRobotControl], developed by the authors, to simulate the virtual instrument, and compute the desired robot torques due to the interface components. This library is used both for the tuning and implementation of the controller. During the simulation stage, the controller is interfaced with a simulated Franka-Emika robot. At implementation stage, the simulated robot is replaced by the real robot, via rerouting of the communication. The virtual model controller runs on a linux PC with the real-time kernel patch and transmits the desired joint torques to the robot at a rate of $1\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$, which are then realized by the robot’s internal joint torque controllers.
TABLE V: Extended scenario tuned parameters (rounded).
    Final
Description     $\bm{\theta}$}   $\hat{\bm{\theta}}$
@@ VII Experimental validation
[…]
The experimental errors of the tip and RCM shown in Fig. 19(e) and Fig. 19(f) vary depending on pose but generally remain well below 10 mm. Their statistics are summarised in Table VI. Although this performance is not accurate enough in terms of absolute placement of the surgical tool tip, we should recall that the tool reference is governed by a surgeon who can adapt the reference accordingly to visual feedback and insights from experience. In that sense, we would argue that an interpretable and compliant response of the robot is more important than absolute accuracy. In general, reference tracking can be improved by increasing the stiffness of the virtual mechanism, if the bandwidth of the actuators allows for it (this is a limitation of the nonsurgical manipulator used in the experiments). Alternatively, performance can be improved by introducing integral action within a slow outer feedback loop, as shown in [Larby2025], or by considering friction correction terms, if a precise model can be estimated. These additional control terms do not necessarily preserve passivity but are entirely compatible with the virtual mechanism approach.
Because of our approach and the structure of the controller, particularly its passivity, we expect that the interconnection of the virtual model controller and the real robot should remain stable. This is despite mismatches between the robot model used for the tuning and the robot itself, such as uncertainty in the inertia parameters and the lack of any joint-friction model. Due to the constraint of passing through the rubber aperture of the surgical phantom, all of the 4 poses tested, represented in Fig. 19(c), deviate from the training pose. This results in a different behaviour to that experienced in simulation and tuning. However, due to the structure of the controller, we still expect reasonable performances.
Notably, if the sensing-communication-actuation loop accurately renders the desired torques, the controller’s passivity guarantees that the closed-loop system will be stable. This stability holds even when the real robot’s dynamics do not match the simulated dynamics and under interaction with external passive systems, such as the instrument interacting with the phantom. However, a loss of performance is expected due the mismatch between ideal and real parameters of the robot.
@@
Figure 19: Experimental performance of the controller while passing through 4 different ’incisions’ on a laparoscopy training phantom, labelled RCM-A through D. Video frames showing RCM-A at t=0s, 1s, …5s, from the side A) and top B) of the phantom respectively. C) Initial poses for of the robot/virtual instrument. D) Tip error x-y trajectories. E) Tip error magnitude. F) RCM error magnitudes. All positions are computed using the kinematics of the robot and measured joint angles.
TABLE VI: RCM and Tip errors mean $\pm$ standard deviation.
Test RCM Error (mm) Tip Error (mm)
RCM-A 1.92 $\pm$ 1.10 4.52 $\pm$ 3.77
RCM-B 2.34 $\pm$ 1.43 5.70 $\pm$ 3.69
RCM-C 2.27 $\pm$ 0.79 5.39 $\pm$ 3.65
RCM-D 2.15 $\pm$ 0.99 5.30 $\pm$ 3.65
VIII Conclusions